At&t is the dominant firm in the local telecommunication industry, which has a total market demand given by Q = 100 - 2P. AT&T has competition from a fringe of four small firms that produce where their individual marginal costs equal the market price. The fringe firms each have total costs given by TCi = 10 Qi + Qi^2. If AT&T's total costs are given by TCa = 10 + 10 Qa, how much does the industry as a whole produce at this price?

Answers:

60

40

80

20

100
Answer #1

Option 1) 60

od = 100-2P, OS Q Res Mga = lo 10000VA, Ti = 109 + 9 i ² Mei= 10+2 q i q= (P-10)/2 then supply of fringe fimos , total, 95 =